# %TODO␣FINAL:␣Replace␣with␣your␣institution␣list.www.retrocausal.ai

### v6

@@ 3.4 Multi-Modality Fusion
Differences with CASA [43]. There are several differences between our method and CASA [43], including: 1) Our method simply requires 2D skeletons (vs. 3D skeletons in CASA). 2) Our method performs self-attention both in spatial and temporal domains (vs. no spatial self-attention in CASA). 3) Our simple heatmap augmentations are based on 2D skeletons (vs. 3D skeletons in CASA). 4) Our method obtains both higher accuracy and better robustness than CASA. 5) Our multi-modality fusion further improves the performance (vs. no multi-modality fusion in CASA).
@@ 4 Experiments
Datasets. We evaluate on IKEA ASM [4], H2O [42], and Penn Action [81]. Following [30, 50, 43], we experiment on the Kallax Drawer Shelf videos of IKEA ASM [4], showing a furniture assembly activity. H2O [42] is a hand-object interaction dataset. We evaluate on the Pouring Milk activity of H2O, following [43]. Penn Action [81] contains sporting activities. We use the same 13 actions of Penn Action as [23, 30, 50, 43]. These datasets are captured with diverse camera viewpoints, actors, backgrounds, and actions, as illustrated in Fig. 5. We use full-body poses for Penn Action [81] and IKEA ASM [4], hand poses for H2O [42], and the same training/testing splits as [23, 30, 50, 43].
Implementation Details. Our approach is implemented in pyTorch [58]. We randomly initialize our model and use ADAM [40] optimization. For Penn Action [81], one model is trained separately for each of the 13 actions. For datasets with a single action such as IKEA ASM [4] and H2O [42], a single model is trained. Due to space limits, more details are available in the supplementary material.
Competing Methods. We test our approach, namely LA2DS (short for Learning by Aligning 2D Skeletons), against state-of-the-art self-supervised temporal video alignment methods, including SAL [55], TCN [62], TCC [23], LAV [30], VAVA [50], and CASA [43]. Moreover, we combine CASA [43] with the graph convolutional network STGCN [78], and integrate LAV [30] with the video transformer ViViT [2] (pre-trained on Kinetics-400 [36]) as additional competing methods.
Evaluation Metrics. We compute four metrics on the validation set, i.e., Phase Classification Accuracy, Phase Progression, Kendall’s Tau, and Average Precision. Please refer to [23, 30, 50, 43] for the definitions. We train the network on the training set and freeze it, and then train an SVM classifier or a linear regressor on top of the frozen features.
@@ 4.1.3 Robustness against Missing Keypoints
H2O CASA [43] $\blacktriangle$ 68.78 66.68 64.80 60.93
@@
To assess the robustness of our method, we perform an experiment where a certain part of the skeleton is removed. Specifically, we randomly remove one finger with a probability $p$ for each frame in the H2O dataset during testing (without any finetuning) and measure the impact of this perturbation on the classification accuracy. The results in Tab. 3 show that LA2DS exhibits a high level of robustness to missing keypoints. For example, when one finger is dropped per frame (i.e., $p=100\%$), the decrease in accuracy for LA2DS is minimal (about $1\%$), whereas for CASA [43], the drop is around $8\%$. This is likely because our method represents each joint by a Gaussian distribution and performs self-attention in the spatial domain to leverage dependencies between joints in the same skeleton, whereas CASA models each joint merely by its coordinates and does not conduct spatial self-attention.
@@ 4.1.4 Robustness against Noisy Keypoints
H2O CASA [43] $\blacktriangle$ 68.78 65.15 65.09 63.69
@@
To examine the generalization of our approach to noisy keypoints, we train and test on skeletons of different qualities. In particular, we use ground truth keypoints provided by the H2O dataset, namely clean keypoints, as well as add random noise to them, yielding noisy keypoints. We measure the impact of this perturbation on the classification accuracy. The results in Tab. 4 show that LA2DS exhibits a high level of robustness against noisy keypoints. For instance, when noisy keypoints are used for both training and testing, the accuracy of CASA [43] is decreased by more than 5%, whereas our method has a drop of less than 2%. Similar to the prior section, our method is more robust to noisy keypoints than CASA, thanks to our heatmap representation and self-attention in the spatial domain.
@@ 4.2.1 Phase Classification Results
Penn Action SAL [55] $\square$ 79.94 81.11 81.79
TCN [62] $\square$ 81.99 82.64 82.78
TCC [23] $\square$ 79.72 81.12 81.35
LAV [30] $\square$ 83.56 83.95 84.25
ViViT [2]+LAV [30] $\square$ 70.66 72.92 74.76
VAVA [50] $\square$ 83.89 84.23 84.48
CASA [43] $\blacktriangle$ 88.55 91.87 92.20
STGCN [78]+CASA [43] $\blacktriangle$ 87.45 90.55 91.37
@@
IKEA ASM SAL [55] $\square$ 21.68 21.72 22.14
TCN [62] $\square$ 25.17 25.70 26.80
TCC [23] $\square$ 24.74 25.22 26.46
LAV [30] $\square$ 29.78 29.85 30.43
VAVA [50] $\square$ 31.66 33.79 32.91
CASA [43] $\blacktriangle$ 21.32 31.52 31.06
@@
H2O TCC [23] $\square$ 43.30 52.48 52.78
LAV [30] $\square$ 35.38 51.66 53.43
CASA [43] $\blacktriangle$ 43.50 62.51 68.78
@@
We first evaluate the effectiveness of our learned representations for action phase classification. Tab. 5 reports the quantitative results of all methods on Penn Action, IKEA ASM, and H2O. It is evident that our method achieves the best overall performance, despite diverse camera viewpoints, actors, backgrounds, and actions (e.g., see Fig. 5). The inferior performance of ViViT [2]+LAV [30] is likely because ViViT [2] has roughly one fifth number of parameters of ResNet-50 [31] used in the original LAV [30]. In contrast, we empirically observe that STGCN [78]+CASA [43] often suffers from overfitting due to low-dimensional input vectors. Also, we notice that RGB-based methods (i.e., LAV, VAVA) often outperform both 2D and 3D skeleton-based methods (i.e., CASA, LA2DS) in limited label settings on IKEA ASM, since different actions may have similar skeletons on IKEA ASM (e.g., picking up front panel, picking up back panel) and context details from RGB data are important for classifying those actions. We will later address this limitation by multi-modality fusion (see Sec. 4.3). The results in Tab. 5 indicate that LA2DS is effective in learning spatiotemporal and contextual features, yielding superior classification performance.
@@ 4.2.2 Phase Progression and Kendall’s Tau Results
Penn Action SAL [55] $\square$ 0.6960 0.7612
TCN [62] $\square$ 0.7217 0.8120
TCC [23] $\square$ 0.6638 0.7012
LAV [30] $\square$ 0.6613 0.8047
ViViT [2]+LAV [30] $\square$ 0.7681 0.7059
VAVA [50] $\square$ 0.7091 0.8053
CASA [43] $\blacktriangle$ 0.9449 0.9728
STGCN [78]+CASA [43] $\blacktriangle$ 0.9061 0.9156
@@
H2O LAV [30] $\square$ 0.5913 0.5323
CASA [43] $\blacktriangle$ 0.9107 0.9438
@@
Tab. 6 presents the phase progression and Kendall’s Tau results on Penn Action and H2O (IKEA ASM is not included due to its repetitive labels [30]). Phase progression measures how well an action progresses over the video, whereas Kendall’s Tau measures how well two sequences are aligned temporally. From Tab. 6, it can be seen that our approach obtains the best performance in Kendall’s Tau, while having comparably best numbers along with CASA [43] in phase progression. This suggests that our method is capable of capturing dependencies between skeletons across sequences, which play a crucial role for temporal alignment.
@@ 4.2.3 Fine-Grained Frame Retrieval Results
Penn Action SAL [55] $\square$ 76.04 75.77 75.61
TCN [62] $\square$ 77.84 77.51 77.28
TCC [23] $\square$ 76.74 76.27 75.88
LAV [30] $\square$ 79.13 78.98 78.90
ViViT [2]+LAV [30] $\square$ 87.58 83.07 80.16
VAVA [50] $\square$ 81.52 80.47 80.67
CASA [43] $\blacktriangle$ 89.90 89.44 89.07
STGCN [78]+CASA [43] $\blacktriangle$ 90.32 89.56 88.41
@@
IKEA ASM SAL [55] $\square$ 15.15 14.90 14.72
TCN [62] $\square$ 19.15 19.19 19.33
TCC [23] $\square$ 19.80 19.64 19.68
LAV [30] $\square$ 23.89 23.65 23.56
VAVA [50] $\square$ 29.58 28.74 28.48
CASA [43] $\blacktriangle$ 28.92 28.88 28.61
@@
H2O LAV [30] $\square$ 47.55 45.56 44.61
CASA [43] $\blacktriangle$ 60.13 59.44 59.01
@@
Here, we examine the performance of our learned representations for fine-grained frame retrieval on Penn Action, IKEA ASM, and H2O, and compare it with previous methods as illustrated in Tab. 7. Following [30], our experimental setup involves using one video from the validation set as the query video, while the remaining videos in the validation set serve as the support set. To retrieve $K$ most similar frames in the support set for each query frame in the query video, we find $K$ nearest neighbors in the embedding space. We then calculate Average Precision at $K$, which represents the percentage of the $K$ retrieved frames with the same action phase label as the query frame. It is obvious from Tab. 7 that our method achieves the best results, outperforming all competing methods across all datasets and all metrics, despite diverse camera viewpoints, actors, backgrounds, and actions (e.g., see Fig. 5). The results in Tab. 7 validate the effectiveness of our learned representations, which capture dependencies between joints in the same skeleton as well as dependencies between skeletons in the same sequence, for fine-grained frame retrieval.
@@ 4.3 Fusion Results
Penn Action LAV [30] $\square$ 84.25 0.6613 0.8047 78.90
VAVA [50] $\square$ 84.48 0.7091 0.8053 80.67
CASA [43] $\blacktriangle$ 92.20 0.9449 0.9728 89.07
CASA [43] $\blacktriangle$ + $\square$ 92.90 0.9494 0.9808 90.25
@@
Tab. 8 presents the multi-modality fusion results on Penn Action only. The complete results on all metrics and datasets are included in the supplementary material. For comparison purposes, we extend CASA [43] by adding branches with RGB inputs and ViViT encoders and performing RGB+3D skeleton fusion via late fusion (similar to our RGB+2D skeleton fusion in Sec. 3.4), yielding the CASA-based fusion baseline. It is evident from Tab. 8 that our fusion model achieves the best results, outperforming all the single-modality methods and the CASA-based fusion baseline and establishing the state-of-the-art. Both our fusion model and the CASA-based fusion baseline outperform their single-modality counterparts respectively, which confirms the benefit of multi-modality inputs for fine-grained human activity understanding tasks.
@@ 4.4 Comparisons with Recent RGB-Based Methods
We compare our approach with recent RGB-based methods, namely VSP [80] (video inputs) and STEPs [63] (video and optical flow inputs). These methods do not employ temporal alignment losses. The results on IKEA ASM are shown in Tab. 9. Our method outperforms VSP and STEPs on very few metrics, i.e., better retrieval results but worse classification results than VSP, and better Acc@0.1 but worse Acc@0.5 and Acc@1.0 than STEPs. Our worse performance is likely due to our compact inputs (e.g., missing context details). Nevertheless, the key advantage of our method is privacy preserving, which is important for several applications involving human action understanding.
@@
IKEA ASM VSP [80] $\square$ 45.00 45.97 47.52 30.23 - -
STEPs [63] $\approx$ + $\square$ 28.59 36.25 37.02 - - -
@@
Supplementary Material. We provide additional results and discussions in the supplementary material. They include our results with different method parts, which analyze sources of performance gains, our results with OpenPose [8] 2D poses, which are better than our above results with projected 2D poses, and discussions of limitations.

### v7

@@ 3.4 Multi-Modality Fusion
Differences with CASA [kwon2022context]. There are several differences between our method and CASA [kwon2022context], including: 1) Our method simply requires 2D skeletons (vs. 3D skeletons in CASA). 2) Our method performs self-attention both in spatial and temporal domains (vs. no spatial self-attention in CASA). 3) Our simple heatmap augmentations are based on 2D skeletons (vs. 3D skeletons in CASA). 4) Our method obtains both higher accuracy and better robustness than CASA. 5) Our multi-modality fusion further improves the performance (vs. no multi-modality fusion in CASA).
@@ 4 Experiments
Datasets. We evaluate on IKEA ASM [ikeaasm], H2O [h2odataset], and Penn Action [pennaction]. Following [haresh2021learning, liu2022learning, kwon2022context], we experiment on the Kallax Drawer Shelf videos of IKEA ASM [ikeaasm], showing a furniture assembly activity. H2O [h2odataset] is a hand-object interaction dataset. We evaluate on the Pouring Milk activity of H2O, following [kwon2022context]. Penn Action [pennaction] contains sporting activities. We use the same 13 actions of Penn Action as [dwibedi2019tcc, haresh2021learning, liu2022learning, kwon2022context]. These datasets are captured with diverse camera viewpoints, actors, backgrounds, and actions, as illustrated in Fig. 5. We use full-body poses for Penn Action [pennaction] and IKEA ASM [ikeaasm], hand poses for H2O [h2odataset], and the same training/testing splits as [dwibedi2019tcc, haresh2021learning, liu2022learning, kwon2022context].
Implementation Details. Our approach is implemented in pyTorch [paszke2017automatic]. We randomly initialize our model and use ADAM [kingma2014adam] optimization. For Penn Action [pennaction], one model is trained separately for each of the 13 actions. For datasets with a single action such as IKEA ASM [ikeaasm] and H2O [h2odataset], a single model is trained. Due to space limits, more details are available in the supplementary material.
Competing Methods. We test our approach, namely LA2DS (short for Learning by Aligning 2D Skeletons), against state-of-the-art self-supervised temporal video alignment methods, including SAL [ishan2016sal], TCN [sermanet2018tcn], TCC [dwibedi2019tcc], LAV [haresh2021learning], VAVA [liu2022learning], and CASA [kwon2022context]. Moreover, we combine CASA [kwon2022context] with the graph convolutional network STGCN [yan2018spatial], and integrate LAV [haresh2021learning] with the video transformer ViViT [arnab2021vivit] (pre-trained on Kinetics-400 [kinetics400]) as additional competing methods.
Evaluation Metrics. We compute four metrics on the validation set, i.e., Phase Classification Accuracy, Phase Progression, Kendall’s Tau, and Average Precision. Please refer to [dwibedi2019tcc, haresh2021learning, liu2022learning, kwon2022context] for the definitions. We train the network on the training set and freeze it, and then train an SVM classifier or a linear regressor on top of the frozen features.
@@ 4.1.3 Robustness against Missing Keypoints
H2O CASA [kwon2022context] $\blacktriangle$ 68.78 66.68 64.80 60.93
@@
To assess the robustness of our method, we perform an experiment where a certain part of the skeleton is removed. Specifically, we randomly remove one finger with a probability $p$ for each frame in the H2O dataset during testing (without any finetuning) and measure the impact of this perturbation on the classification accuracy. The results in Tab. 3 show that LA2DS exhibits a high level of robustness to missing keypoints. For example, when one finger is dropped per frame (i.e., $p=100\%$), the decrease in accuracy for LA2DS is minimal (about $1\%$), whereas for CASA [kwon2022context], the drop is around $8\%$. This is likely because our method represents each joint by a Gaussian distribution and performs self-attention in the spatial domain to leverage dependencies between joints in the same skeleton, whereas CASA models each joint merely by its coordinates and does not conduct spatial self-attention.
@@ 4.1.4 Robustness against Noisy Keypoints
H2O CASA [kwon2022context] $\blacktriangle$ 68.78 65.15 65.09 63.69
@@
To examine the generalization of our approach to noisy keypoints, we train and test on skeletons of different qualities. In particular, we use ground truth keypoints provided by the H2O dataset, namely clean keypoints, as well as add random noise to them, yielding noisy keypoints. We measure the impact of this perturbation on the classification accuracy. The results in Tab. 4 show that LA2DS exhibits a high level of robustness against noisy keypoints. For instance, when noisy keypoints are used for both training and testing, the accuracy of CASA [kwon2022context] is decreased by more than 5%, whereas our method has a drop of less than 2%. Similar to the prior section, our method is more robust to noisy keypoints than CASA, thanks to our heatmap representation and self-attention in the spatial domain.
@@ 4.2.1 Phase Classification Results
Penn Action SAL [ishan2016sal] $\square$ 79.94 81.11 81.79
TCN [sermanet2018tcn] $\square$ 81.99 82.64 82.78
TCC [dwibedi2019tcc] $\square$ 79.72 81.12 81.35
LAV [haresh2021learning] $\square$ 83.56 83.95 84.25
ViViT [arnab2021vivit]+LAV [haresh2021learning] $\square$ 70.66 72.92 74.76
VAVA [liu2022learning] $\square$ 83.89 84.23 84.48
CASA [kwon2022context] $\blacktriangle$ 88.55 91.87 92.20
STGCN [yan2018spatial]+CASA [kwon2022context] $\blacktriangle$ 87.45 90.55 91.37
@@
IKEA ASM SAL [ishan2016sal] $\square$ 21.68 21.72 22.14
TCN [sermanet2018tcn] $\square$ 25.17 25.70 26.80
TCC [dwibedi2019tcc] $\square$ 24.74 25.22 26.46
LAV [haresh2021learning] $\square$ 29.78 29.85 30.43
VAVA [liu2022learning] $\square$ 31.66 33.79 32.91
CASA [kwon2022context] $\blacktriangle$ 21.32 31.52 31.06
@@
H2O TCC [dwibedi2019tcc] $\square$ 43.30 52.48 52.78
LAV [haresh2021learning] $\square$ 35.38 51.66 53.43
CASA [kwon2022context] $\blacktriangle$ 43.50 62.51 68.78
@@
We first evaluate the effectiveness of our learned representations for action phase classification. Tab. 5 reports the quantitative results of all methods on Penn Action, IKEA ASM, and H2O. It is evident that our method achieves the best overall performance, despite diverse camera viewpoints, actors, backgrounds, and actions (e.g., see Fig. 5). The inferior performance of ViViT [arnab2021vivit]+LAV [haresh2021learning] is likely because ViViT [arnab2021vivit] has roughly one fifth number of parameters of ResNet-50 [he2016deep] used in the original LAV [haresh2021learning]. In contrast, we empirically observe that STGCN [yan2018spatial]+CASA [kwon2022context] often suffers from overfitting due to low-dimensional input vectors. Also, we notice that RGB-based methods (i.e., LAV, VAVA) often outperform both 2D and 3D skeleton-based methods (i.e., CASA, LA2DS) in limited label settings on IKEA ASM, since different actions may have similar skeletons on IKEA ASM (e.g., picking up front panel, picking up back panel) and context details from RGB data are important for classifying those actions. We will later address this limitation by multi-modality fusion (see Sec. 4.3). The results in Tab. 5 indicate that LA2DS is effective in learning spatiotemporal and contextual features, yielding superior classification performance.
@@ 4.2.2 Phase Progression and Kendall’s Tau Results
Penn Action SAL [ishan2016sal] $\square$ 0.6960 0.7612
TCN [sermanet2018tcn] $\square$ 0.7217 0.8120
TCC [dwibedi2019tcc] $\square$ 0.6638 0.7012
LAV [haresh2021learning] $\square$ 0.6613 0.8047
ViViT [arnab2021vivit]+LAV [haresh2021learning] $\square$ 0.7681 0.7059
VAVA [liu2022learning] $\square$ 0.7091 0.8053
CASA [kwon2022context] $\blacktriangle$ 0.9449 0.9728
STGCN [yan2018spatial]+CASA [kwon2022context] $\blacktriangle$ 0.9061 0.9156
@@
H2O LAV [haresh2021learning] $\square$ 0.5913 0.5323
CASA [kwon2022context] $\blacktriangle$ 0.9107 0.9438
@@
Tab. 6 presents the phase progression and Kendall’s Tau results on Penn Action and H2O (IKEA ASM is not included due to its repetitive labels [haresh2021learning]). Phase progression measures how well an action progresses over the video, whereas Kendall’s Tau measures how well two sequences are aligned temporally. From Tab. 6, it can be seen that our approach obtains the best performance in Kendall’s Tau, while having comparably best numbers along with CASA [kwon2022context] in phase progression. This suggests that our method is capable of capturing dependencies between skeletons across sequences, which play a crucial role for temporal alignment.
@@ 4.2.3 Fine-Grained Frame Retrieval Results
Penn Action SAL [ishan2016sal] $\square$ 76.04 75.77 75.61
TCN [sermanet2018tcn] $\square$ 77.84 77.51 77.28
TCC [dwibedi2019tcc] $\square$ 76.74 76.27 75.88
LAV [haresh2021learning] $\square$ 79.13 78.98 78.90
ViViT [arnab2021vivit]+LAV [haresh2021learning] $\square$ 87.58 83.07 80.16
VAVA [liu2022learning] $\square$ 81.52 80.47 80.67
CASA [kwon2022context] $\blacktriangle$ 89.90 89.44 89.07
STGCN [yan2018spatial]+CASA [kwon2022context] $\blacktriangle$ 90.32 89.56 88.41
@@
IKEA ASM SAL [ishan2016sal] $\square$ 15.15 14.90 14.72
TCN [sermanet2018tcn] $\square$ 19.15 19.19 19.33
TCC [dwibedi2019tcc] $\square$ 19.80 19.64 19.68
LAV [haresh2021learning] $\square$ 23.89 23.65 23.56
VAVA [liu2022learning] $\square$ 29.58 28.74 28.48
CASA [kwon2022context] $\blacktriangle$ 28.92 28.88 28.61
@@
H2O LAV [haresh2021learning] $\square$ 47.55 45.56 44.61
CASA [kwon2022context] $\blacktriangle$ 60.13 59.44 59.01
@@
Here, we examine the performance of our learned representations for fine-grained frame retrieval on Penn Action, IKEA ASM, and H2O, and compare it with previous methods as illustrated in Tab. 7. Following [haresh2021learning], our experimental setup involves using one video from the validation set as the query video, while the remaining videos in the validation set serve as the support set. To retrieve $K$ most similar frames in the support set for each query frame in the query video, we find $K$ nearest neighbors in the embedding space. We then calculate Average Precision at $K$, which represents the percentage of the $K$ retrieved frames with the same action phase label as the query frame. It is obvious from Tab. 7 that our method achieves the best results, outperforming all competing methods across all datasets and all metrics, despite diverse camera viewpoints, actors, backgrounds, and actions (e.g., see Fig. 5). The results in Tab. 7 validate the effectiveness of our learned representations, which capture dependencies between joints in the same skeleton as well as dependencies between skeletons in the same sequence, for fine-grained frame retrieval.
@@ 4.3 Fusion Results
Penn Action LAV [haresh2021learning] $\square$ 84.25 0.6613 0.8047 78.90
VAVA [liu2022learning] $\square$ 84.48 0.7091 0.8053 80.67
CASA [kwon2022context] $\blacktriangle$ 92.20 0.9449 0.9728 89.07
CASA [kwon2022context] $\blacktriangle$ + $\square$ 92.90 0.9494 0.9808 90.25
@@
Tab. 8 presents the multi-modality fusion results on Penn Action only. The complete results on all metrics and datasets are included in the supplementary material. For comparison purposes, we extend CASA [kwon2022context] by adding branches with RGB inputs and ViViT encoders and performing RGB+3D skeleton fusion via late fusion (similar to our RGB+2D skeleton fusion in Sec. 3.4), yielding the CASA-based fusion baseline. It is evident from Tab. 8 that our fusion model achieves the best results, outperforming all the single-modality methods and the CASA-based fusion baseline and establishing the state-of-the-art. Both our fusion model and the CASA-based fusion baseline outperform their single-modality counterparts respectively, which confirms the benefit of multi-modality inputs for fine-grained human activity understanding tasks.
@@ 4.4 Comparisons with Recent RGB-Based Methods
We compare our approach with recent RGB-based methods, namely VSP [zhang2023modeling] (video inputs) and STEPs [shah2023steps] (video and optical flow inputs). These methods do not employ temporal alignment losses. The results on IKEA ASM are shown in Tab. 9. Our method outperforms VSP and STEPs on very few metrics, i.e., better retrieval results but worse classification results than VSP, and better Acc@0.1 but worse Acc@0.5 and Acc@1.0 than STEPs. Our worse performance is likely due to our compact inputs (e.g., missing context details). Nevertheless, the key advantage of our method is privacy preserving, which is important for several applications involving human action understanding.
@@
IKEA ASM VSP [zhang2023modeling] $\square$ 45.00 45.97 47.52 30.23 - -
STEPs [shah2023steps] $\approx$ + $\square$ 28.59 36.25 37.02 - - -
@@
Supplementary Material. We provide additional results and discussions in the supplementary material. They include our results with different method parts, which analyze sources of performance gains, our results with OpenPose [cao2019openpose] 2D poses, which are better than our above results with projected 2D poses, and discussions of limitations.